Lambda Corporation has current liabilities of $432,000, a quick ratio of 1.9, inventory turnover of 4.6, and a current ratio of 3.3. What is the cost of goods sold for Lambda Corporation?
Answer #1

Lambda Corporation has current liabilities of $432,000, a quick ratio of 1.9, inventory turnover of 4.6, and a current ratio of 3.3. What is the cost of goods sold for Lambda Corporation?

1)

Current Ratio = 3.3

Current Asset/current liabilities = 3.3

Current Asset = 3.3*432000

Current Asset = 1425600

2)

Quick Ratio = 1.9

Quick Asset/current liabilities = 1.90

Quick Asset = 1.90*432000

Quick Asset = 820800

3)

Quick Asset = Current asset - Inventory

Inventory = Current asset - Quick Asset

Inventory = 1425600-820800

Inventory = 604800

4)

Inventory turnover = cost of goods sold / Inventory

4.6 = cost of goods sold/ 604800

cost of goods sold = 4.6*604800

cost of goods sold = $ 2,782,080

Answer

cost of goods sold for Lambda Corporation = $ 2,782,080